Thermal Rising, Inc., makes paragliders for sale through specialty sporting goods stores. The company has a standard paraglider model, but also makes custom-designed paragliders. Management has designed an activity-based costing system with the following activity cost pools and activity rates:

Activity Cost Pool Activity Rate
Supporting direct labor $ 20 per direct labor-hour
Order processing $ 198 per order
Custom design processing $ 266 per custom design
Customer service $ 430 per customer

Management would like an analysis of the profitability of a particular customer, Big Sky Outfitters, which has ordered the following products over the last 12 months:

Standard
Model
Custom
Design
Number of gliders 15 2
Number of orders 2 2
Number of custom designs 0 2
Direct labor-hours per glider 27.50 33.00
Selling price per glider $ 1,625 $ 2,380
Direct materials cost per glider $ 444 $ 574

The company’s direct labor rate is $18 per hour.

Required:

Using the company’s activity-based costing system, compute the customer margin of Big Sky Outfitters. (Round your intermediate calculations and final answer to the nearest whole dollar amount. Loss amounts should be entered with a minus sign.)
Answer #1
Standard Custom Big Sky Outfitters
Sales Revenue        24,375        4,760                   29,135
Less: Costs
Direct Material           6,660        1,148                      7,808
Direct labor           7,425        1,188                      8,613
Overheads
Supporting Direct labor           8,250        1,320                      9,570
Order Processing              396            396                         792
Custom Design                 -              532                         532
Customer Service                         430
Total cost                   27,745
Customer Margin                      1,390

А в с 1 Standard Custom Big Sky Outfitters 2 Sales Revenue =15*1625 =2*2380 =B2+C2 3 Less: Costs 4. Direct Material =15*444 =